Equation is :

Mg(s)+2HCl(aq)→MgCl2(aq)+H2(g)

Temperature: 21 ℃

Mass of magnesium metal used: 0.058 g

Volume and molarity of HCl used: 10 mL

Volume of hydrogen gas obtained: 70.5 mL



- Find these theoretical calculations:
What is the atom economy (%)?
What is the percent of reactant atoms incorporated into desired product, H2 (%)?
- Find these experimental calculations?
What is the percent yield (%)?
Reaction mass efficiency (%)?


I have my own answers but for the percent yield I found a result of 120 % which is I know impossible.
So it is really frustrating, because I have no idea what my mistake is.
Thanks a lot.
1
Expert's answer
2020-01-29T04:43:45-0500

Atom economy (atom efficiency/percentage) is the conversion efficiency of a chemical process in terms of all atoms involved and the desired products produced.

Atom economy: 2 (H2)=24(Mg)+73(2HCl)=2/97=0.0206 or 2.06%


The product yield is a percentage of the theoretically possible yield according to the reaction equation.

Percent yield (H2)=70.5mL/54mL=1.305 or 130.5%, 70,5 mL -

actual volume, 54 mL- theoretical volume. 54mL=22.4 l/mole*0.002 mole, 0.002=0.058g/24 (Mg)


Mass reaction efficiency is defined as the ratio of the mass obtained product to the mass of the purified product, divided by the sum of the masses of all substances introduced intothe reaction mixture during the experiment/